The world’s largest gold bar in the toi gold museum in japan can be seen and touched by visitors. it has an irregular shape with dimensions about 17.9 in. by 8.9 in. by 6.7 inches, with a volume of 12.94 l. determine the mass and weight (1 kg = 2.205 lb).


1
Expert's answer
2022-02-14T17:10:03-0500

Most of the information here is not required for solution (dimensions, conversion factor). Knowing that the density of gold is 19.32 kg/L,


"m=\rho\times{V}=19.32\frac{kg}{L}\times12.94\ L=250.0\ kg"

"W=m\times{g}=250.0\ kg\times9.8\frac{N}{kg}=2450\ N"


Answer: mass = 250.0 kg; weight = 2450 N